OpenStudy (anonymous):

WILL MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!! J is the tallest on his basketball team. He is 1 1/2 times as tall as the shortest girl in 6th grade, who is 4 1/4 feet tall, how tall is J?

OpenStudy (anonymous):

8 1/2?

OpenStudy (anonymous):

Actually I think thats wrong. All the calculators i have used got 6 and 3/8. And i got the same answer. but i have to stay true to my title...

OpenStudy (anonymous):

\[1\dfrac{1}{2}\times4\dfrac{1}{4} = 6\dfrac{3}{8}\]

OpenStudy (anonymous):

that's correct answer